# Quasi-exactly solvable potentials in Wigner-Dunkl quantum mechanics

### 09 Jan 2024 v1

@@ 1 Introduction
In 1950, Wigner introduced the use of the reflection operator in quantum mechanics [1] and Yang then applied it to the harmonic oscillator [2]. Sets of differential-difference operators associated with finite reflection groups were independently introduced by Dunkl [3]. These operators, now referred to as Dunkl operators, have proved very useful for studying polynomials in several variables with discrete symmetry [4]. In physics too, such operators have been employed for bosonizing supersymmetric quantum mechanics [5] or generalizations thereof [6] and have provided a tool for proving the superintegrability of some models [7].
More recently, deformations of the momentum operator of such a type in quantum mechanics have attracted much attention and have given rise to what is often called Wigner-Dunkl quantum mechanics. For instance, the Dunkl oscillator and the Dunkl Coulomb problems in the plane were investigated [8, 9], as well as their generalizations to three dimensions [10, 11]. Coherent states [12], a generalization of shape invariance in supersymmetric quantum mechanics [13], and relativistic effects [14, 15] have also been considered. The possibility of using Dunkl derivatives with two or three parameters has been investigated [16, 17]. Very recently, some rationally-extended potentials, whose wavefunctions can be expressed in terms of exceptional orthogonal polynomials have been studied in the Wigner-Dunkl formalism [18, 19].
@@ 3 QES oscillator potential in the plane
where $a>0$, $n\in\mathbb{N}$, and $n+1$ known solutions for $k=0,1,\ldots,n$.
@@ 4 QES Coulomb potential in the plane
In polar coordinates, the Dunkl Coulomb potential in the plane [9] can be treated in a way rather similar to that of the oscillator, the potential $\frac{1}{2}\rho^{2}$ being replace by $-\frac{\alpha}{2\rho}$ (with $\alpha>0$) in (18) and (19). [^2] The potential $\alpha/\rho$ of [9] is replaced here by $-\alpha/(2\rho)$ to conform with the choice made in [21]. The counterpart of (19) looks like the radial equation of the (three-dimensional) Coulomb problem with Hamiltonian
@@
${\cal R}^{(a,b,1)}_{\begin{smallmatrix}0\\ 1\end{smallmatrix},\nu}(\rho)\propto e^{-\frac{1}{2}a\rho^{2}-b\rho}\rho^{2\nu}\left(2a\rho+b\pm\sqrt{b^{2}+2a(4\nu+2\mu_{1}+\mu_{2}+1)}\right),$ (29)
@@ 5 Conclusion
Second, the Dunkl isotropic oscillator and the Dunkl Coulomb potentials in the plane have been generalized by using the similarity between their radial equation and that of the standard isotropic oscillator or standard Coulomb potential, respectively. As a result, we have obtained for $n=0,1,2,\ldots$, QES Dunkl oscillator-like potentials with $n+1$ known eigenstates, as well as sets of $n+1$ QES Dunkl Coulomb-like potentials with a given known energy.

### 19 Mar 2024 v2

@@ 1 Introduction
In 1950, Wigner considered a deformed version of the Heisenberg algebra in quantum mechanics [1] and then Yang introduced the reflection operator to discuss the harmonic oscillator problem along these lines [2]. Sets of differential-difference operators associated with finite reflection groups were independently introduced by Dunkl [3]. These operators, now referred to as Dunkl operators, have proved very useful for studying polynomials in several variables with discrete symmetry [4]. In physics too, such operators have been employed for bosonizing supersymmetric quantum mechanics [5] or generalizations thereof [6] and have provided a tool for proving the superintegrability of some models [7].
More recently, deformations of the momentum operator of such a type in quantum mechanics have attracted much attention and have given rise to what is often called Wigner-Dunkl quantum mechanics. For instance, the Dunkl oscillator and the Dunkl-Coulomb problems in the plane were investigated [8, 9], as well as their generalizations to three dimensions [10, 11]. Coherent states [12], a generalization of shape invariance in supersymmetric quantum mechanics [13], and relativistic effects [14, 15] have also been considered. The possibility of using Dunkl derivatives with two or three parameters has been investigated [16, 17]. Very recently, some rationally-extended potentials, whose wavefunctions can be expressed in terms of exceptional orthogonal polynomials have been studied in the Wigner-Dunkl formalism [18, 19].
@@ 3 QES oscillator potential in the plane
where $a>0$, $n\in\mathbb{N}$, and with $n+1$ known solutions for $k=0,1,\ldots,n$.
@@ 4 QES Coulomb potential in the plane
In polar coordinates, the Dunkl-Coulomb potential in the plane [9] can be treated in a way rather similar to that of the oscillator, the potential $\frac{1}{2}\rho^{2}$ being replace by $-\frac{\alpha}{2\rho}$ (with $\alpha>0$) in (18) and (19). [^2] The potential $\alpha/\rho$ of [9] is replaced here by $-\alpha/(2\rho)$ to conform with the choice made in [21]. The counterpart of (19) looks like the radial equation of the (three-dimensional) Coulomb problem with Hamiltonian
@@
${\cal R}^{(a,b,1)}_{\begin{smallmatrix}0\\ 1\end{smallmatrix},\nu}(\rho)\propto e^{-\frac{1}{2}a\rho^{2}-b\rho}\rho^{2\nu}\left(2a\rho+b\pm\sqrt{b^{2}+2a(4\nu+2\mu_{1}+2\mu_{2}+1)}\right),$ (29)
@@ 5 Conclusion
Second, the Dunkl isotropic oscillator and the Dunkl-Coulomb potentials in the plane have been generalized by using the similarity between their radial equation and that of the standard isotropic oscillator or standard Coulomb potential, respectively. As a result, we have obtained for $n=0,1,2,\ldots$, QES Dunkl oscillator-like potentials with $n+1$ known eigenstates, as well as sets of $n+1$ QES Dunkl Coulomb-like potentials with a given known energy.